Unit Information with BWIP, Weighted Average Method Jackson Products produces a barbeque sauce using three departments: Cooking, Mixing, and Bottling. In the Cooking Department, all materials are added at the beginning of the process. Output is measured in ounces. The production data for July are as follows: Production: Units in process, July 1, 70% complete* 10,800 Units completed and transferred out 88,000 Units in process, July 31, 80% complete* 15,900 * With respect to conversion costs. Required: 1. Prepare a physical flow schedule for July. Jackson Products Physical Flow Schedule For the Month of July Units to account for: Total units to account for Units accounted for: Units completed and transferred out: Total units accounted for 2. Prepare an equivalent units schedule for July using the weighted average method. Enter percentages as whole numbers. Jackson Products Equivalent Units Schedule Direct Materials Conversion Costs Units completed Add: Units in ending work in process × Percentage complete: x % x % Equivalent units of output 3. What if you were asked to calculate the FIFO units beginning with the weighted average equivalent units? Calculate the weighted average equivalent units by subtracting out the prior-period output found in BWIP. FIFO Equivalent Units Direct Materials Conversion Costs
Answer #1

1)

Jackson Products
Physical Flow Schedule
For the Month of July
Units to account for:
Units, beginning work in process (70% complete) 10800
Units started during July (103900-10800) 93100
Total units to account for 103900
Units accounted for:
Units completed and transferred out:
From beginning work in process 10800
Started and completed 77200 88000
Units in ending work in process (80% complete) 15900
Total units accounted for 103900

Units started and completed= Units completed and transferred out-From beginning work in process

= 88000-10800= 77200

2)

Jackson Products
Equivalent Units Schedule
Direct Materials Conversion Costs
Units completed 88000 88000
Add: Units in ending work in process*Percentage complete:
15900*100% 15900 -
15900*80% - 12720
Equivalent units of output 103900 100720

3)

Direct materials Conversion costs
Equivalent units of output using weighted average method 103900 100720
Less: Beginning work in process (10800*100%)= 10800 (10800*70%)= 7560
Equivalent units of output using FIFO method 93100 93160